A health insurer has determined that the reasonable and customary fee for a certain medical procedure is $1220. They suspect that the average fee charged by one particular clinic for this procedure is higher than $1220. The insurer wants to perform a hypothesis test to determine whether their suspicions is correct. Preliminary data analysis indicate that it is reasonable to apply a z-test. Assume that :sd =$238, n=22, and the significance level is 0.10. Express the decision criterion for the hypothesis test in terms of x-bar.
Answer #1

H0:Null Hypothesis: \mu = 1220

HA: Alternative Hypothesis: \mu > 1220 (Claim)

SE = s/n

= 238/V22

= 50.7418

Test Statistic is given by:

Z = (\bar{x} - 1220)/50.7418
\alpha = 0.10

One Tailed - Right Side

From Table, critical value of Z= 1.28

Decision criterion for the hypothesis test is given by:

Reject H0 if

Z > 1.28

i.e.,

(\bar{x} - 1220)/50.7418 > 1.28

i.e.,

\bar{x}- 1220 > 64.9495

i.e.,

\bar{x} > 1284.9495

So,

Decision criterion for the hypothesis test in terms of \bar{x} is given by:

Reject H0 if

\bar{x} > 1284.9495